Novak Company purchased equipment for $230,000 on October 1, 2020. It is estimated that the equipment will have a useful life of 8 years and a salvage value of $14,000. Estimated production is 36,000 units and estimated working hours are 20,000. During 2020, Novak uses the equipment for 520 hours and the equipment produces 1,000 units. Compute depreciation expense under each of the following methods. (Novak is on a calendar-year basis ending December 31.)

a. Straight line method for 2020

b. Activity Method (units of output) for 2020

c. Activity Method (working hours) for 2020

d. Sum-of-the-years'-digits method for 2022

e. Double-declining balance method for 2021
Answer #1

A. Straight line method.

>> Depreciation per annum = ( $ 230,000 - $ 14,000 ) / 8 = $ 27,000

>> Depreciation for 2020 = $ 27,000 * ( 3 /12 ) = $ 6,750.

________________________________________________________________________________________

B. Units of output.

>> Depreciation per unit = ( $ 230,000 - $ 14,000 ) / 36,000 = $ 6

>> Depreciation for 2020 = 1,000 * $ 6 = $ 6,000.

__________________________________________________________________________________________

C. Working ours.

>> Depreciation per hours = ( $ 230,000 - $ 14,000 ) / 20,000 = $ 10.8

>> Depreciation for 2020 = 520 * $ 10.8 = $ 5,616.

______________________________________________________________________________________________

D. Sum of digits method

>> Sum of years = 8 ( 8 +1 ) / 2 = 36.

>> Depreciation for 2022 = ( $ 230,000 - $ 14,000 ) * ( 6.75 / 36 ) = $ 40,500.

___________________________________________________________________________________________

E . Double declining balance

>> Depreciation rate = 200 / 8 = 25 %.

>> Depreciation for 2020 = $ 230,000 * 25 % * ( 3 /12 ) = $ 14,375.

>> Depreciation for 2021 = ( $ 230,000 - $ 14,375 )* 25 % = $ 53,906